OpenStudy (anonymous):

Another logarithm problem Solve the equation: 256^5x=64^x+4

OpenStudy (vijay):

u mean 256^(5x)=64^(x+4)

OpenStudy (anonymous):

5xlog (256)=x log (64) + log(4) using log with base 2 5x*8 = 6x + 2 34x = 2 x = 17

OpenStudy (anonymous):

x = 1/17 oops

OpenStudy (anonymous):

yes vijay 256^(5x)=64^(x+4)

OpenStudy (anonymous):

i'm still confused on this answer...

OpenStudy (vijay):

x = 12/14